OpenStudy (anonymous):

Aiden takes 0.10 s to throw a baseball, which leaves his hand with a velocity of 40 m/s. The ball's acceleration is _____.

OpenStudy (anonymous):

40/ 0.10 m/s^2

OpenStudy (anonymous):

400 m/s^2

OpenStudy (anonymous):

(Vf-Vi)/t = a (40m/s - 0 m/s)/.1s = a 40 m/s / .1s = a a = 400 m/s^2